Peter is 3 times as old as joseph.in 6 years time joseph's age will be 1/2 of peter's age. Find peter's age in 4 years time
Answer by macston(5194)

P=Peter's age; J=Joseph's age
P=3J 2(J+6)=P+6
3J=2(J+6)-6
3J=2J+12-6
3J=2J+6 subtract 2J from each side
3J-2J=2J+6-2J
J=6 Joseph is 6
P=3J=3(6)=18 Peter is 18
In 4 years Peter will be 22 years old
CHECK
in 6 years Joseph's age will be 1/2 Peter's age
J+6=1/2(P+6)
6+6=1/2(18+6)
12=1/2(24)
12=12